OpenStudy (waheguru):

What is the difference between h=-0.0625d(d-56) and h=-0.0625d(d-56)(d-56)

jimthompson5910 (jim_thompson5910):

the first one will turn into a quadratic while the second one will turn into a cubic, so that's one difference

OpenStudy (waheguru):

I mean in terms of graphing and x-intetrcepts

OpenStudy (waheguru):

Do both of them only have one x-intercept (56,0)

jimthompson5910 (jim_thompson5910):

oh, both will have the same x-intercepts, but the graphs will cross them in a different way

jimthompson5910 (jim_thompson5910):

both have an x-intercept of x = 0 and both will have the graph cross through the x-axis at this point the first will have an x-intercept of x = 56 and it will cross through the second will have an x-intercept of x = 56, but it won't cross through: it will just touch the x-axis and not completely cross over

OpenStudy (waheguru):

So the first equation will have the x-intercepts of (0,0) and (56,0) but then the second one will only have (56,0) as an x-intercept?

jimthompson5910 (jim_thompson5910):

no the second also has an x-intercept of x = 0, sorry i should have said that too

OpenStudy (waheguru):

Then how would one know if there is only one x-intercept in the parabola?

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (agent0smith):

By setting it equal to zero: \[\large 0=-0.0625d(d-56)(d-56)\] Use the zero product property, if \[\large (a) (b) (c) = 0\] then either a=0, b=0 or c=0.

OpenStudy (waheguru):

Say the equation only had one x-intercept how would you show it?

OpenStudy (agent0smith):

and \[\large 0 =-0.0625d(d-56) \]

OpenStudy (agent0smith):

You'd show it by finding them from what i just showed.

OpenStudy (waheguru):

I still don't understand how we can can show that the equation only has one x-intercept>

OpenStudy (agent0smith):

From my above post, 0=−0.0625d(d−56)(d−56) Use the zero product property, if (a)(b)(c)=0 then either a=0, b=0 or c=0. This is how you show it.

OpenStudy (agent0smith):

0=(−0.0625d)(d−56)(d−56) maybe that will help.

OpenStudy (waheguru):

So that has only 1 x-intercept right but then this h=-0.0625d(d-56 Has 2? right...?

OpenStudy (waheguru):

Has 2 x-intercepts ...

OpenStudy (agent0smith):

No, it does not have only one intercept. 0=(−0.0625d)(d−56)(d−56) Use the zero product property, if (a)(b)(c)=0 then either a=0, b=0 or c=0.

OpenStudy (agent0smith):

Set them all equal to zero, and find x.

OpenStudy (waheguru):

but the equation is u=a(x-s)(x-r) there is no A B or C. Im sorrry but I have not learned the zero product property

OpenStudy (waheguru):

y=a(x-s)(x-r)

OpenStudy (waheguru):

or no B and C?

OpenStudy (agent0smith):

if 0=a(x-s)(x-r) then either a=0, x-s=0 or x-r=0

OpenStudy (waheguru):

Yes I get that, say there was a parabola that faced down and had a step pattern of 1,3,5 and the vertex was at (5,0) how would I make this into an equation in the factored form

OpenStudy (waheguru):

anyone?

OpenStudy (waheguru):

@agent0smith

OpenStudy (agent0smith):

What's a step pattern? http://www.mesacc.edu/~scotz47781/mat120/notes/graph_quads/vertex_form/graph_quads_vertex_form.html ^ vertex form